OpenStudy (anonymous):

x + x + x = x3 A. Always B. Sometimes C. Never i am so confused

jimthompson5910 (jim_thompson5910):

if that x3 is x times 3, then it would be the same as 3x however, it looks like they meant to say x^3 or \(\large x^3\)

jimthompson5910 (jim_thompson5910):

that is NOT true because it *should* be x*x*x = x^3 or \[\Large x*x*x = x^3\]

jimthompson5910 (jim_thompson5910):

and \[\Large x + x + x = 3x\]

OpenStudy (anonymous):

GOD BLESS JIM Lol but thank you for explaining this

jimthompson5910 (jim_thompson5910):

you're welcome